# Nonlinear morphoelastic energy based theory for stimuli responsive elastic shells

### v2

@@ 3.1 Expansion of the invariants of $\bf{C}_{e}$
Then we rewrite the in-plane contribution of the determinant in terms of the fundamental forms as (see Supplementary Material Supplementary Material for detailed algebra of the derivation)
@@
where we introduced the mean curvature $H$ and Gaussian curvature $K$ in current configuration and the mean curvature $\bar{H}$ and Gaussian curvature $\bar{K}$ in the intermediate configuration. Finally, by taking their ratio and expanding the expression in the functions $\phi$ and $\bar{\phi}$ we get the desired form. Naturally we obtain
@@
To derive this form, we first recall that, in our setting, the second invariant can be rewritten as Supplementary Material
@@
The form of the function $\phi$ is motivated by the choice in Ozenda and Virga [Ozenda2021] and the remarks therein presented are valid also in our framework, with the addition of considering the presence of a natural curvature and a non-elastic stimulus. The polynomial expansion for $\bar{\phi}$ is justified by similarity with the choice for $\phi$.
@@ 4 Constitutive Modeling and Reduced Energies
The general kinematic description presented in Section 3 must be complemented with constitutive relationships specific to the problem under investigation. In what follows, we consider two representative materials: an incompressible Neo-Hookean material and a compressible Ciarlet–Geymonat model. These choices not only reflect widely used forms in the literature, but also enable validation of our derivation by recovering the known reduced model for non-morphoelastic plates from Ozenda and Virga [Ozenda2021].
@@ 4.1 Incompressible Neo-Hookean Energy
For an incompressible material, the coefficients of the through-the-thickness function in eq. (42) can be derived solely from kinematic considerations [Ozenda2021]. Incompressibility implies $J=1$ and we also limit ourselves to the case where $J_{s}=1$, i.e., the mid-surface is inextensible. Inserting eq. (41) and eq. (42) into eq. (3.1) yields the following conditions for the coefficients in the through-the-thickness function $\phi$ as
@@
Upon integration over the thickness, the energy functional in eq. (5) becomes:
$\Psi_{nc,2d}=\int_{\mathcal{S}_{0}}\left[h_{0}\bar{\rho}_{1}\left(\frac{\mu^{*}}{2}w^{\textup{NH}}_{s}+p(J_{s}-1)\right)+\frac{h_{0}^{3}}{12}\frac{1}{2\bar{\rho}_{1}^{3}}\left(\mu^{*}w^{\textup{NH}}_{b}+2\bar{\rho}_{1}^{3}pf_{p}\right)\right]\sqrt{\det(\bar{\mathrm{a}}_{\alpha\beta})}\,\mathrm{d}\eta_{1}\mathrm{d}\eta_{2}+O(h^{5})\,$, (46)
where the integral is evaluated with respect to the coordinates in the reference configuration. For compactness, we present the quantities in the reduced functional when setting $\bar{\rho}_{2}=\bar{\rho}_{3}=0$. We refer to the Supplementary Material Supplementary Material for their form including all the terms of the through-the-thickness stimulus. The stretching and bending terms are given by:
$w^{\textup{NH}}_{s}=$ $\mathrm{a}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}-2,$ (47)
$w^{\textup{NH}}_{b}=$ $\bar{\rho}_{1}^{3}f_{1}w_{s}^{NH}+2\bar{\rho}_{1}^{4}f_{2}\mathrm{b}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}+2\bar{\rho}_{1}^{4}f_{3}\mathrm{a}_{\alpha\beta}\bar{\mathrm{b}}^{\alpha\beta}$
$+\bar{\rho}_{1}^{6}\left(-4\mathrm{b}_{\alpha\beta}\bar{\mathrm{b}}^{\alpha\beta}+\mathrm{c}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}+3\mathrm{a}_{\alpha\beta}\bar{\mathrm{c}}^{\alpha\beta}\right)+\bar{\rho}_{1}^{2}f_{4}\,$, (48)
@@
$f_{1}$ $=\bar{\rho}_{1}^{3}\bar{K},$ (49)
$f_{2}$ $=2\bar{\rho}_{1}^{2}\bar{H}-\bar{\rho}_{1},$ (50)
$f_{3}$ $=-2\bar{\rho}_{1}^{2}\bar{H},$ (51)
$f_{4}$ $=2\bar{\rho}_{1}^{4}\left(6H(H-\bar{H})-K+\bar{K}\right)-8\bar{\rho}_{1}^{3}\bar{H}+4\bar{\rho}_{1}^{2},$ (52)
$f_{p}$ $=\bar{\rho}_{1}^{3}(J_{s}-1)\bar{K}+6\bar{\rho}_{1}\left(J_{s}(H-\bar{H})H\bar{\rho}_{1}^{2}-\bar{\rho}_{1}HJ_{s}\right)\,.$ (53)
@@ 4.2 Compressible Ciarlet–Geymonat Energy
$\left(\mu^{*}\right)^{-1}\Psi_{cg}=\int_{\mathcal{S}_{0}}\left[h_{0}w^{CG}_{s}+\frac{h_{0}^{3}}{12}w^{CG}_{b}\right]\sqrt{\det(\bar{\mathrm{a}}_{\alpha\beta})}\,\mathrm{d}\eta_{1}\mathrm{d}\eta_{2}+O(h^{5})\,$, (55)
and we introduce the rescaled bulk modulus as $\lambda=\lambda^{*}/\mu^{*}$. For compactness, here we assume again that $\bar{\rho}_{2}=\bar{\rho}_{3}=0$ and we refer to the Supplementary Material Supplementary Material for the full form of the reduced energy. The stretching contribution reads
@@
Finally, it can be shown that, as in the non-morphoelastic plate case [Ozenda2021], the function $\rho_{3}$ does not contribute to the energy up to $O(h_{0}^{5})$ and thus remains undetermined.
@@ 4.3 Linearization and bending-stretching coupling
It is illustrative to understand the behavior of our model against a linear elastic shell model, based on the Koiter energy with distortions, as used in [Pezzulla2018PRL]. The Koiter shell model relies on a series of assumptions. Beyond the fact that its validity is limited to small strains, (i) it requires the use, in this exact order [Efrati2009], of the plane stress hypothesis to rewrite the constitutive relationship and the plane strain hypothesis to constrain the through-the-thickness behavior (when it is derived from a 3D setting [Steigmann2013]) and (ii) in the energy formulation, stretching and bending strains are decoupled, with the former entering only at order $O(h_{0})$ and the latter only at order $O(h_{0}^{3})$.
@@
where we use the shape operator $\bar{S}^{\alpha}_{\beta}=\bar{\mathrm{a}}^{\alpha\gamma}\bar{\mathrm{b}}_{\gamma\beta}$.
This linearization reveals two aspects of the behavior of our model in the presence of small strains and moderate rotations. If the undeformed configuration is flat, bending and stretching are fully decoupled, although with different coefficients due to the lack of plane stress enforcement (for comparison, the reader can check the formulation used, for example, in Pezzulla et al. [Pezzulla2018PRL]); if the undeformed configuration is curved, the term $w_{\textup{gnl}}$ is non-zero and the non-linearities due to the interaction with the geometry are retained [Steigmann2013, Wood2019].
@@ 5.1 Minimization of the energy functional
The energy functional derived in Section 4.2 is implemented in COMSOL Multiphysics, following the same procedure presented in Pezzulla et al. [Pezzulla2018PRL]. Both the stretching (the term proportional to $h_{0}$) and bending (the term proportional to $h_{0}^{3}$) components in the energy are functions of the first, second, and third fundamental forms. In turn, the fundamental forms can be expressed as functions of the embeddings of the undeformed and deformed mid-surfaces, namely $\mathbf{X}$ and $\mathbf{x}$. As a result, the total elastic energy of the shell becomes a second-order functional of $\mathbf{x}$, which we minimize in COMSOL Multiphysics.
@@ 5.2 Eversion of a cap
Pezzulla et al. [Pezzulla2018PRL] investigated the eversion of a spherical cap, whose mid-surface in the deformed configuration is parametrized as
@@
This stimulus imposes a change the curvature of size $\kappa$, the only controlling parameter, without affecting the area of the mid-surface fixed, and without adding any contribution through-the-thickness. This stimulus also implies $\bar{\mathrm{c}}_{\alpha\beta}=2\bar{H}\bar{\mathrm{b}}_{\alpha\beta}-\bar{K}\bar{\mathrm{a}}_{\alpha\beta}$ where $\bar{K}=\overset{\circ}{K}+\kappa^{2}-2\kappa R_{0}^{-1}$ and $\bar{H}=\overset{\circ}{H}+\kappa$, from straightforward algebra. The energy is minimized by fixing to zero the displacement of the apex of the cap at $s=0$ and tracking the vertical component $w$ of the displacement of the free end of the cap at $s=\Theta$ .
@@
Figure 2: Snapping of a spherical cap. Natural curvature at snapping, normalized by the theoretical prediction from [Pezzulla2018PRL], as a function of the vertical displacement at the free end $w$, normalized by the radius of the shell $R_{0}$, for $\Theta=\pi/6$ and $\lambda=0.01,0.1,1,10$ when (a) $\rho_{1}=1$ and $\rho_{2}=0$ and (b) $\rho_{1}$ and $\rho_{2}$ fixed by eqs. (57) and (108).
Fig. 2 shows the curvature stimulus at snap-through $\kappa_{s}$, normalized by the theoretical prediction from Pezzulla et al. [Pezzulla2018PRL] obtained by using the classic morphoelastic Koiter shell theory [Efrati2009], function of the normalized displacement at the apex, $w/R_{0}$. In fig. 2(a) we see that snap-through (indicated by the presence of a maximum in the curves) is present any value of $\lambda$ but it decreases when moving towards the limit of incompressibility, $\lambda\rightarrow\infty$, suggesting the possibility for snapping to be eventually suppressed. Experiments [Pezzulla2018PRL] however, shows that snap-through is expected for morphing spherical caps also of incompressible (or nearly incompressible) materials. If we relax the plane strain requirement, thus allowing $\rho_{1}$ and $\rho_{2}$ to be fixed by eqs. (57) and (108), fig. 2(b) shows how snap-through happens for any value of the shear to bulk moduli ratio and it remains close to the experimental values.
@@
Although the morphoelastic Koiter theory has been proved to correctly model the snap-through events in spherical caps [Pezzulla2018PRL], the combination of an appropriate linearized theory that retains geometric non-linearities (in case of small strains and moderate rotations our model reduced to the form shown in Section 4.3) and the appropriate modeling of the through-the-thickness behavior may be the correct approach to be used to unveil the fundamental aspects in the physics of eversion morphing, also allowing the correct modelling of plane-strain and plane-stress scenarios.
Figure 3: Dimensionless imposed curvature at snapping, $\kappa_{s}R_{0}$ in terms of the rescaled shallowness $\theta\sqrt{R_{0}/h_{0}}$. The solid line represents the theoretical prediction from [Pezzulla2018PRL], valid for an incompressible Koiter shell model. The colored curves represent the result of our model for different values of $\lambda=0.01,0.1,1,10$ when $\rho_{1}$ and $\rho_{2}$ are fixed by eqs. (57) and (108).

### v3

@@ 3.1 Expansion of the invariants of $\bf{C}_{e}$
Then we rewrite the in-plane contribution of the determinant in terms of the fundamental and target forms as (see appendix A.1 for detailed algebra of the derivation)
@@
where we introduced the mean curvature $H$ and Gaussian curvature $K$ in current configuration and the target mean curvature $\bar{H}$ and target Gaussian curvature $\bar{K}$ in the intermediate configuration. The quantities $\bar{H}$ and $\bar{K}$ are computed as
$\bar{H}=\frac{1}{2}\mathrm{tr}\left(\bar{\mathrm{b}}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}\right),\quad\bar{K}=\frac{\det(\bar{\mathrm{b}}_{\alpha\beta})}{\det(\bar{\mathrm{a}}_{\alpha\beta})}\,.$ (36)
Although these definitions resemble algebraically the ones used to compute mean and Gaussian curvatures, we must not intend $\bar{H}$ and $\bar{K}$ as curvatures associated with an embedding. Finally, by taking their ratio and expanding the expression in the functions $\phi$ and $\bar{\phi}$ we get the desired form. Naturally we obtain
@@
To derive this form, we first recall that, in our setting, the second invariant can be rewritten as shown in appendix A.2
@@
The form of the function $\phi$ is motivated by the choice in Ozenda and Virga [24] and the remarks therein presented are valid also in our framework, with the addition of considering the presence of a natural curvature and a non-elastic stimulus. The polynomial expansion for $\bar{\phi}$ is justified by similarity with the choice for $\phi$.
@@ 4 Constitutive Modeling and Reduced Energies
The general kinematic description presented in Section 3 must be complemented with constitutive relationships specific to the problem under investigation. In what follows, we consider two representative materials: an incompressible Neo-Hookean material and a compressible Ciarlet–Geymonat model. These choices not only reflect widely used forms in the literature, but also enable validation of our derivation by recovering the known reduced model for non-morphoelastic plates from Ozenda and Virga [24].
@@ 4.1 Incompressible Neo-Hookean Energy
For an incompressible material, the coefficients of the through-the-thickness function in eq. (43) can be derived from kinematic considerations [24]. Incompressibility implies $J=1$ and we first limit ourselves to the case where the mid-surface is inextensible, i.e. $J_{s}=1$, and we derive an energy functional where this latter constraint is enforced variationally. Inserting eq. (42) and eq. (43) into eq. (3.1) yields the following conditions for the coefficients in the through-the-thickness function $\phi$ as
@@
Upon integration over the thickness, also expanding the pressure field as $p=p_{0}(\eta^{\alpha})Z+p_{1}(\eta^{\alpha})Z+O(Z^{2})$ (adding the term proportional to $Z^{2}$ would only introduce a redundant constraint at the truncation order used here), the energy functional in eq. (5) becomes:
$\Psi_{nc,2d}=\int_{\mathcal{S}_{0}}\left[h_{0}w^{\textup{NH}}_{s}+\frac{h_{0}^{3}}{12}w^{\textup{NH}}_{b}\right]\sqrt{\det(\bar{\mathrm{a}}_{\alpha\beta})}\,\mathrm{d}\eta^{1}\mathrm{d}\eta^{2}+O(h^{5})\,$, (47)
where the integral is evaluated with respect to the coordinates in the reference configuration. For compactness, we present the quantities in the reduced functional when setting $\bar{\rho}_{2}=\bar{\rho}_{3}=0$. We refer to the appendix B.1 for their form including all the terms of the through-the-thickness stimulus. The stretching and bending terms, terms proportional to $h_{0}$ and $h_{0}^{3}$ respectively, are given by
$w^{\textup{NH}}_{s}=$ $\bar{\rho}_{1}\left(\frac{\mu^{*}}{2}\left(\mathrm{a}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}-2\right)+p_{0}(J_{s}-1)\right),$ (48)
$w^{\textup{NH}}_{b}=$ $\frac{\mu^{*}}{2}\bar{\rho}_{1}^{3}\left(2\frac{\bar{K}}{\bar{\rho}_{1}}w^{\textup{NH}}_{s}+2\left(H-\bar{H}\right)\mathrm{b}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}-4\bar{H}\mathrm{a}_{\alpha\beta}\bar{\mathrm{b}}^{\alpha\beta}+p_{0}f_{p0}+p_{1}f_{p1}\right$.
$\left.+\left(-4\mathrm{b}_{\alpha\beta}\bar{\mathrm{b}}^{\alpha\beta}+\mathrm{c}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}+3\mathrm{a}_{\alpha\beta}\bar{\mathrm{c}}^{\alpha\beta}\right)+f_{1}\right)\,$, (49)
@@
$f_{1}$ $=\left(4(H-\bar{H})(4H+\bar{H})-2K+2\bar{K}\right),$ (50)
$f_{p0}$ $=\frac{24}{\mu^{*}}HJ_{s}(H-\bar{H})\,$ (51)
$f_{p1}$ $=\frac{4}{\mu^{*}\bar{\rho}_{1}}\left(-2HJ_{s}+\bar{H}J_{s}+\bar{H}\right)\,.$ (52)
@@
It is possible to relax the constraint $J_{s}=1$ and derive an energy functional where the full constraint $J=1$ is imposed kinematically [24]. Following the same procedure employed in the derivation of the expressions in eqs. (44) – (46) , the coefficients in the through-the-thickness function $\phi$ modify as
$\rho_{1}(\eta^{\alpha})$ $=\frac{\bar{\rho}_{1}}{J_{s}},$ (54)
$\rho_{2}(\eta^{\alpha})$ $=\frac{\bar{\rho}_{1}^{2}}{J_{s}^{2}}\left(\bar{H}-H\right)+\frac{\bar{\rho}_{2}}{J_{s}},$ (55)
$\rho_{3}(\eta^{\alpha})$ $=\frac{\bar{\rho}_{3}}{J_{s}}+2\frac{\bar{\rho}_{1}\bar{\rho}_{2}}{J_{s}^{2}}\left(J_{s}\bar{H}-H\right)+\frac{\bar{\rho}_{1}^{3}}{3J_{s}^{3}}\left(6H\left(H-J_{s}\bar{H}\right)-(K-J_{s}^{2}\bar{K})\right)\,.$ (56)
Upon integration over the thickness the terms in the energy in eq. (47) in this case becomes
$w^{\textup{NH}}_{s}=$ $\frac{\bar{\rho}_{1}\mu^{*}}{2}\left(\mathrm{a}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}-3+\frac{1}{J_{s}^{2}}\right),$ (57)
$w^{\textup{NH}}_{b}=$ $\frac{\mu^{*}\bar{\rho}_{1}^{3}}{2J_{s}^{4}}\left(J_{s}^{4}\bar{K}w_{s}^{NH}+2J_{s}\left(H-\bar{H}J_{s}\right)\mathrm{b}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}-4J_{s}^{4}\bar{H}\mathrm{a}_{\alpha\beta}\bar{\mathrm{b}}^{\alpha\beta}\right$.
$\left.+J_{s}^{2}\left(-4J_{s}\mathrm{b}_{\alpha\beta}\bar{\mathrm{b}}^{\alpha\beta}+\mathrm{c}_{\alpha\beta}\bar{\mathrm{a}}^{\alpha\beta}+3J_{s}^{2}\mathrm{a}_{\alpha\beta}\bar{\mathrm{c}}^{\alpha\beta}\right)\right$. (58)
$\left.+2\left(8H^{2}-6H\bar{H}J_{s}-K+J_{s}^{2}\left(-2\bar{H}^{2}+\bar{K}\right)\right)\right)\,$, (59)
valid when imposing $\bar{\rho}_{2}=\bar{\rho}_{3}=0$.
@@ 4.2 Compressible Ciarlet–Geymonat Energy
$\left(\mu^{*}\right)^{-1}\Psi_{cg}=\int_{\mathcal{S}_{0}}\left[h_{0}w^{CG}_{s}+\frac{h_{0}^{3}}{12}w^{CG}_{b}\right]\sqrt{\det(\bar{\mathrm{a}}_{\alpha\beta})}\,\mathrm{d}\eta^{1}\mathrm{d}\eta^{2}+O(h^{5})\,$, (60)
and we introduce the rescaled bulk modulus as $\lambda=\lambda^{*}/\mu^{*}$. For compactness, here we assume again that $\bar{\rho}_{2}=\bar{\rho}_{3}=0$ and we refer to the appendix B.2 for the full form of the reduced energy. The stretching contribution reads
@@
Finally, it can be shown that, as in the non-morphoelastic plate case [24], the function $\rho_{3}$ does not contribute to the energy up to $O(h_{0}^{5})$ and thus remains undetermined.
@@ 4.3 Linearization and bending-stretching coupling
It is illustrative to understand the behavior of our model against a linear elastic shell model, based on the Koiter energy with distortions, as used in [25]. The Koiter shell model relies on a series of assumptions. Beyond the fact that its validity is limited to small strains, (i) it requires the use, in this exact order [8], of the plane stress hypothesis to rewrite the constitutive relationship and the plane strain hypothesis to constrain the through-the-thickness behavior (when it is derived from a 3D setting [31]) and (ii) in the energy formulation, stretching and bending strains are decoupled, with the former entering only at order $O(h_{0})$ and the latter only at order $O(h_{0}^{3})$.
@@
where we introduce the symbol $\bar{S}^{\alpha}_{\beta}=\bar{\mathrm{a}}^{\alpha\gamma}\bar{\mathrm{b}}_{\gamma\beta}$ that identifies a generalized target shape operator.
This linearization reveals two aspects of the behavior of our model in the presence of small strains and moderate rotations. If the undeformed configuration is flat, bending and stretching are fully decoupled, although with different coefficients due to the lack of plane stress enforcement (for comparison, the reader can check the formulation used, for example, in Pezzulla et al. [25]); if the undeformed configuration is curved, the term $w_{\textup{gnl}}$ is non-zero and the non-linearities due to the interaction with the geometry are retained [31, 36].
@@ 5.1 Minimization of the energy functional
The energy functional derived in Section 4.2 is implemented in COMSOL Multiphysics, following the same procedure presented in Pezzulla et al. [25]. Both the stretching (the term proportional to $h_{0}$) and bending (the term proportional to $h_{0}^{3}$) components in the energy are functions of the first, second, and third fundamental forms. In turn, the fundamental forms can be expressed as functions of the embeddings of the undeformed and deformed mid-surfaces, namely $\mathbf{X}$ and $\mathbf{x}$. As a result, the total elastic energy of the shell becomes a second-order functional of $\mathbf{x}$, which we minimize in COMSOL Multiphysics.
@@ 5.2 Eversion of a cap
Pezzulla et al. [25] investigated the eversion of a spherical cap, whose mid-surface in the deformed configuration is parametrized as
@@
This stimulus imposes a change the curvature of size $\kappa$, the only controlling parameter, without affecting the area of the mid-surface fixed, and without adding any contribution through-the-thickness. This stimulus also implies $\bar{\mathrm{c}}_{\alpha\beta}=2\bar{H}\bar{\mathrm{b}}_{\alpha\beta}-\bar{K}\bar{\mathrm{a}}_{\alpha\beta}$ where $\bar{K}=\overset{\circ}{K}+\kappa^{2}-2\kappa R_{0}^{-1}$ and $\bar{H}=\overset{\circ}{H}+\kappa$, from straightforward algebra and consistently with the constraint on $\bar{\mathrm{c}}_{\alpha\beta}$ over which we derive the energy functionals. The energy is minimized by fixing to zero the displacement of the apex of the cap at $s=0$ and tracking the vertical component $w$ of the displacement of the free end of the cap at $s=\Theta$ .
@@
Figure 2: Snapping of a spherical cap. Natural curvature at snapping, normalized by the theoretical prediction from [25], as a function of the vertical displacement at the free end $w$, normalized by the radius of the shell $R_{0}$, for $\Theta=\pi/6$ and $\lambda=0.01,0.1,1,10$ when (a) $\rho_{1}=1$ and $\rho_{2}=0$ and (b) $\rho_{1}$ and $\rho_{2}$ fixed by eqs. (62) and (65).
Fig. 2 shows the curvature stimulus at snap-through $\kappa_{s}$, normalized by the theoretical prediction from Pezzulla et al. [25] obtained by using the classic morphoelastic Koiter shell theory [8], function of the normalized displacement at the apex, $w/R_{0}$. In fig. 2(a) we see that snap-through (indicated by the presence of a maximum in the curves) is present any value of $\lambda$ but it decreases when moving towards the limit of incompressibility, $\lambda\rightarrow\infty$, suggesting the possibility for snapping to be eventually suppressed. Experiments [25] however, shows that snap-through is expected for morphing spherical caps also of incompressible (or nearly incompressible) materials. If we relax the plane strain requirement, thus allowing $\rho_{1}$ and $\rho_{2}$ to be fixed by eqs. (62) and (65), fig. 2(b) shows how snap-through happens for any value of the shear to bulk moduli ratio and it remains close to the experimental values.
@@
Although the morphoelastic Koiter theory has been proved to correctly model the snap-through events in spherical caps [25], the combination of an appropriate linearized theory that retains geometric non-linearities (in case of small strains and moderate rotations our model reduced to the form shown in Section 4.3) and the appropriate modeling of the through-the-thickness behavior may be the correct approach to be used to unveil the fundamental aspects in the physics of eversion morphing, also allowing the correct modelling of plane-strain and plane-stress scenarios.
Figure 3: Dimensionless imposed curvature at snapping, $\kappa_{s}R_{0}$ in terms of the rescaled shallowness $\theta\sqrt{R_{0}/h_{0}}$. The solid line represents the theoretical prediction from [25], valid for an incompressible Koiter shell model. The colored curves represent the result of our model for different values of $\lambda=0.01,0.1,1,10$ when $\rho_{1}$ and $\rho_{2}$ are fixed by eqs. (62) and (65).